OpenStudy (swag):

Fred is fencing his yard, which is shaped as a rhombus. He used 4x meters of fencing material for one side and (x + 9) meters for the adjacent side. What is the perimeter of the yard? 12 meters 36 meters 48 meters 96 meters

OpenStudy (hoblos):

all sides of a rhombus are equal so 4x=x+9 3x=9 x=3 so each side is = 4x=4(3)=12 perimeter 4*side = 48 meters

OpenStudy (swag):

You should be a Ambassador

OpenStudy (hoblos):

:D:D